# Causal Deep Learning

### v2

@@ 1 Introduction
Causal explanations specify the causes, the mechanism, and the conditions for replicating an observed effect [47, 24, 92]. Quantitatively, causality is the direct relationship between two events, A and B, where “A $\xrightarrow[]{\mbox{causes}}$ B” means “the effect of A is B”, a measurable and experimentally repeatable phenomena. Once verified with either experimental or observational studies, the statement “the effect of A is B” stays true regardless of new discoveries and changes in knowledge [31]. Causal explanations are a factual understanding, whereas interpretations– a concept also advocated for the development of trustworthy artificial intelligence – are an understanding relative to a specific reference frame, such as a particular point of view or knowledge base. As new knowledge emerges, prior interpretations may be deemed to be inaccurate or invalid, which can undermine their reliability and usefulness in the development of trustworthy artificial intelligence. Interpretations are open to reinterpretation. Therefore, explanations should not be confused with interpretations or speculations masquerading as explanations [49, 48].
The validity and robustness of causal explanations depend on causal model specifications in conjunction with the experimental designs used for acquiring training data [63]. Generative artificial intelligence research that focuses on unsupervised deep representation learning [21, 7, 54, 71] is not well suited for drawing causal conclusions. (Determining the causal factors from unsupervised training data is an ill-posed inverse problem since different combinations of the same causal factors can result in the same outcome.) This prior unsupervised generative artificial intelligence research briefly addressed the connection to tensor (multilinear) factor analysis, a supervised learning approach that models the causal mechanisms that generates data [82, 85], estimates the effects of causes [84, 89], as well as, the causes of effects given an estimated forward model and constraints on the solution set [86, 78]. Data and its representations – such as those derived from principal or independent component analysis etc.– are factorized into a set of invariant, discriminant, and disentangled causal representations [77][(4.15)][82]. The causal representations maximize the ratio of inter- to intra-class scatter [82]. Thus, a causal tensor model is not only generative, but discriminative. The most successful generative deep neural networks are also supervised with generative and/or discriminative properties [43, 70, 53]. However, their goal has been prediction rather than causal inference.
@@
Forward causal inference is a hypothesis-driven process, as opposed to a data-driven process, that models the mechanism of data formation and estimates the effects of interventions and counterfactuals [58, 36, 68, 80]. Conventional statistics and machine learning, which model data distributions, predict one variable co-observed with another, or perform time series forecasting. Inverse causal “inference” estimates the causes of effects given an estimated forward model and constraints on the solution set [78, 26].
@@ 2.4 Neural Network Architecture
${\mathbf{d}}_{i_{\mbox{\tiny{1}}},\dots,i_{\mbox{\tiny{m}}}\dots i_{\mbox{\tiny{M}}}}$ $=$ ${\mathbf{\mathcal{Z}}}\times_{\mbox{\tiny{1}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{1}}}\times\dots\times_{\mbox{\tiny{m}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\times\dots\times_{\mbox{\tiny{M}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}$ (12)
@@ 2.5 Causal Deep Networks and Scalable Tensor Factor Analysis:
Figure 3: Deep neural network. Subfigures (a-e) depict (3140). (a) The mode matrix computation ${\mathbf{U}}_{\mbox{\tiny{m}}}$ may be thought as a constrained cluster-based PCA that is rewritten in terms of block SVDs. Matrix4izing $\mathbf{\mathcal{D}}$} may be viewed as a concatenation of “cluster” data. The matrix $\mathbf{W}$} transforms the basis matrix ${\mathbf{V}}^{(n)}_{\mbox{\tiny{0}}}$ such that the causal factor representation ${\mathbf{U}}_{\mbox{\tiny{m}}}$ is the same regardless of cluster membership. In a tensor model, there are $M$ different constrained cluster-based PCAs. (b) Mode matrix ${\mathbf{U}}_{\mbox{\tiny{m}}}$ computation using a single autoencoder-decoder. (c) Mode matrix computation as a hierarchy of autoencoder-decoders (i.e., feed-forward [45]), (d) Mode matrix computation written as a deep learning model (e) Concurrent-autoencoders; i.e., constrained cluster-based autoencoders. (f) Forward causal model with a set of capsules implemented by deep neural networks. For a parallel, synchronized or asynchronous computation, we break the chain links and shuttle causal information, ${\mathbf{U}}_{\mbox{\tiny{m}}}$, between capsules to compute ${\mathbf{\mathcal{X}}}(t+1)$ for the next iteration. (g) Each capsule in (f) may be replaced with a part-based deep neural network by permuting the rows in ${{\mathbf{D}}{}^{\hbox{\tiny{T}}}}_{\mbox{\tiny{$[m]$}}}$ with $\mathbf{P}$}, segmented by ${\mathbf{H}}_{\mbox{\tiny{m}}}$, which is efficiently trained with a part-based hierarchy of autoencoders, Fig. 4.
@@
Computing causal factor representations, the mode matrices ${\mathbf{U}}_{\mbox{\tiny{m}}}$ of an MPCA tensor model, is equivalent to computing a $M$ different of mutually constrained, cluster-based PCA [77, pg.38-40], Fig. 3a. When dealing with data that can be separated into clusters, the standard machine learning approach is to compute a separate PCA. When data from different clusters are generated by the same underlying process (e.g., facial images of the same people under different viewing conditions), the data blocks can be concatenated in the measurement mode and the common causal factor can be modeled by one PCA. However, for a scalable solution, we employ block algebra (31) and compute a set of constrained cluster-based PCAs, i.e., a set of concurrent PCAs. Thus, we define a constrained, cluster-based PCA as the computation of a set of PCA basis vectors, such that the latent representation is constrained to be the invariant of the cluster membership.
@@ 3 Inverse Causal Question: “Why?”
Multilinear tensor factor analysis constrains causal factor representations to be unitary vectors. Multilinear projection [86, 78] relies on this constraint and performs multiple regularized regressions. One or more unlabeled test observations that are not part of the training data set are simultaneously projected into the causal factor spaces,

### v3

@@ 1 Introduction
Causal explanations specify the causes, the mechanism, and the conditions for replicating an observed effect [47, 24, 92]. Quantitatively, causality is the direct relationship between two events, A and B, where “A $\xrightarrow[]{\mbox{causes}}$ B” means “the effect of A is B”, a measurable and experimentally repeatable phenomena. Once verified with either experimental or observational studies, the statement “the effect of A is B” stays true regardless of new discoveries and changes in knowledge [31]. Causal explanations are a factual understanding, whereas interpretations – a concept also advocated for the development of trustworthy artificial intelligence [48] – are an understanding relative to a specific reference frame, such as a particular point of view or knowledge base. As new knowledge emerges, prior interpretations may be deemed to be inaccurate or invalid, which can undermine their reliability and usefulness in the development of trustworthy artificial intelligence. Interpretations are open to reinterpretation. Therefore, explanations should not be confused with interpretations or speculations masquerading as explanations [49, 48].
The validity and robustness of causal explanations depend on causal model specifications in conjunction with the experimental designs used for acquiring training data [63]. Generative artificial intelligence research that focuses on unsupervised deep neural networksis not well suited for drawing causal conclusions. (Determining the causal factors from unsupervised training data is an ill-posed inverse problem since different combinations of the same causal factors can result in the same outcome.) Prior unsupervised deep neural network research [21, 7, 54, 71] briefly addressed the connection to tensor (multilinear) factor analysis. Tensor (multilinear) factor analysis is a supervised learning approach that models the causal mechanisms that generate data [82, 85], estimates the effects of causes [84, 89], as well as the causes of effects, given an estimated forward model and constraints on the solution set [86, 78]. Data and its representations – such as those derived from principal or independent component analysis etc.– are factorized into a set of invariant, discriminant, and disentangled causal representations [77][(4.15)][82]. These causal representations maximize the ratio of inter- to intra-class scatter [82]. Thus, a causal tensor model is not only generative, but discriminative. In fact, the most successful deep neural network approaches are also supervised with generative and/or discriminative properties [43, 70, 53]. However, their goal has been prediction rather than causal inference.
@@
Forward causal inference is a hypothesis-driven process, as opposed to a data-driven process, that models the mechanism of data formation and estimates the effects of interventions or counterfactuals [58, 36, 68, 80]. Inverse causal “inference” estimates the causes of effects given an estimated forward model and constraints on the solution set [78, 26]. By comparison, conventional statistics and machine learning model data distributions, predict a variable co-observed with another, or perform time series forecasting.
@@ 2.4 Neural Network Architecture
${\mathbf{d}}_{i_{\mbox{\tiny{1}}},\dots,i_{\mbox{\tiny{m}}}\dots i_{\mbox{\tiny{M}}}}$ $=$ ${\mathbf{\mathcal{T}}}\times_{\mbox{\tiny{1}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{1}}}\times\dots\times_{\mbox{\tiny{m}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{m}}}\times\dots\times_{\mbox{\tiny{M}}}\hat{{\mathbf{u}}}{}^{\hbox{\tiny{T}}}_{\mbox{\tiny{M}}}$ (12)
@@ 2.5 Causal Deep Networks and Scalable Tensor Factor Analysis:
Figure 3: Deep neural network. Subfigures (a-e) depict (3140) and  [77, pg.38-40].
(a) The mode matrix computation ${\mathbf{U}}_{\mbox{\tiny{m}}}$ may be thought as a constrained cluster-based PCA that is rewritten in terms of block SVDs. Matrix4izing $\mathbf{\mathcal{D}}$} may be viewed as a concatenation of “cluster” data. The matrix $\mathbf{W}$} transforms the basis matrix ${\mathbf{V}}^{(n)}_{\mbox{\tiny{0}}}$ such that the causal factor representation ${\mathbf{U}}_{\mbox{\tiny{m}}}$ is the same regardless of cluster membership. In a tensor model, there are $M$ different constrained cluster-based PCAs. (b) Mode matrix ${\mathbf{U}}_{\mbox{\tiny{m}}}$ computation using a single autoencoder-decoder. (c) Mode matrix computation as a hierarchy of autoencoder-decoders (i.e., feed-forward [45]), (d) Mode matrix computation written as a deep learning model (e) Concurrent-autoencoders; i.e., constrained cluster-based autoencoders. (f) Forward causal model with a set of capsules implemented by deep neural networks. For a parallel, synchronized or asynchronous computation, we break the chain links and shuttle causal information, ${\mathbf{U}}_{\mbox{\tiny{m}}}$, between capsules to compute ${\mathbf{\mathcal{X}}}(t+1)$ for the next iteration. (g) Each capsule in (f) may be replaced with a part-based deep neural network by permuting the rows in ${{\mathbf{D}}{}^{\hbox{\tiny{T}}}}_{\mbox{\tiny{$[m]$}}}$ with $\mathbf{P}$}, segmented by ${\mathbf{H}}_{\mbox{\tiny{m}}}$, which is efficiently trained with a part-based hierarchy of autoencoders, Fig. 4.
@@
Computing causal factor representations, the mode matrices ${\mathbf{U}}_{\mbox{\tiny{m}}}$ of an MPCA tensor model, is equivalent to computing a $M$ different of mutually constrained, cluster-based PCA,  [77, pg.38-40], Fig. 3a. When dealing with data that can be separated into clusters, the standard machine learning approach is to compute a separate PCA. When data from different clusters are generated by the same underlying process (e.g., facial images of the same people under different viewing conditions), the data blocks can be concatenated in the measurement mode and the common causal factor can be modeled by one PCA. However, for a scalable solution, we employ block algebra (31) and compute a set of constrained cluster-based PCAs, i.e., a set of concurrent PCAs. Thus, we define a constrained, cluster-based PCA as the computation of a set of PCA basis vectors, such that the latent representation is constrained to be the invariant of the cluster membership.
@@ 3 Inverse Causal Question: “Why?”
Multilinear tensor factor analysis constrains causal factor representations to be unitary vectors. Multilinear projection [86, 78] relies on this constraint and performs multiple regularized regressions. One or more unlabeled test observations that are not part of the training data set are multilinearly projected into the causal factor spaces,